# Investigating Context-Faithfulness in Large Language Models: The Roles of Memory Strength and Evidence Style

### v1

@@ 5 Conclusion
We investigate how context-faithful LLMs are to external evidence across two datasets, PopQA and NQ datasets, using LLaMA2.7B, LLaMA2.70B, ChatGPT, and GPT-4. Our findings highlight the critical role of memory strength in shaping LLM behavior. There is a clear positive correlation between memory strength and memory answer ratio. Furthermore, we demonstrate that paraphrasing significantly enhances the context-faithfulness of LLMs across various models and datasets. In conclusion, our results emphasize the influence of memory strength on the context-faithfulness of LLMs and show that paraphrasing direct evidence improves the receptiveness of LLMs to external evidence. These findings offer valuable insights for advancing research in retrieval-augmented generation and context-based LLM applications.
@@
Appendix A Additional Study
A.1 Order of Options
Figure 4: Impact of Option orders on Memory and Counter Ratios Across NQ and popQA Datasets. Either the memory answer ("mem first") or the counter answer ("ctr first") is introduced first to four models.
Dataset Evidence Style S # LLaMA2.7B LLaMA2.70B ChatGPT GPT-4
$R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$
MA first
NQ Direct 1 7.26 92.75 0.0 3.11 96.89 0.0 18.54 76.38 5.08 48.51 49.48 2.01
Direct + Paraphrase 3 3.26 96.74 0.0 1.11 98.71 0.18 8.99 87.59 3.43 11.34 87.21 1.46
CMA first
Direct 1 22.26 77.73 0.0 19.13 80.38 0.5 34.48 61.82 3.71 49.19 47.99 2.82
Direct + Parapharse 3 4.8 95.11 0.1 8.72 90.39 0.89 18.63 78.96 2.41 17.76 80.02 2.22
MA first
popQA Direct 1 0.7 99.3 0.0 1.31 98.49 0.2 3.51 94.58 1.91 12.99 84.79 2.22
Direct + Paraphrase 3 0.3 99.7 0.0 0.51 99.29 0.2 1.32 98.38 0.3 1.32 98.48 0.2
CMA first
Direct 1 59.37 40.41 0.22 5.75 93.27 0.98 6.22 91.96 1.82 21.44 75.24 3.32
Direct + Paraphrase 3 16.78 83.22 0.0 1.74 98.05 0.22 1.82 97.75 0.43 2.79 96.78 0.43
Table 3: Results of LLM Receptiveness to Different Evidence Styles Across NQ and popQA Datasets. The table presents the MA ratio ($R_{m}$), CMA ratio ($R_{c}$), and UCT ratio ($R_{u}$) for Direct Evidence and Direct + Paraphrase Evidence with CMA first and MA first scenarios. All the ratios are in %.
Across all four models (LLaMA2.7B, LLaMA2.70B, ChatGPT, and GPT-4), we observe a consistent trend: MA ratio ($R_{m}$) under “CMA first” is significantly higher than that under “MA first”. Evaluations under “CMA first” demonstrate that LLMs are less context-faithful.
To further demonstrate the effect of the order of options on $R_{m}$, we compare the performance of experiments with “CMA first” and “MA first” under two evidence styles: direct evidence with one sentence and direct + paraphrase with three sentences. The results are presented in Table 3. The results show that, for different evidence styles, $R_{m}$ is consistently higher in the “CMA first” compared to the “MA first”. Comparing the results under the “CMA first”, the $R_{m}$ of direct + paraphrase with three sentences is significantly lower than that with direct evidence with one sentence. This demonstrates that paraphrasing direct evidence is an effective method for decreasing $R_{m}$. Our conclusion remains unchanged.
A.2 Case Study
The higher $R_{m}$ of LLaMA2.7B may be attributed to its weakness of reasoning ability. From Table 2, we observe that on the NQ dataset, the $R_{m}$ of LLaMA2.7B is higher than that of LLaMA2.70B. This phenomenon is counter-intuitive, as a weaker memory strength in LLMs corresponds to a lower $R_{m}$ for the other three LLMs (LLaMA2.70B, ChatGPT, and GPT-4). To investigate the cause of this issue, we prompt LLaMA2.7B to provide a rationale alongside the answer. We discover that LLaMA2.7B seems to exhibit reasoning errors. Below is an example that shows this phenomenon.
Question: who has the most big ten championships in football
Options:
C: Uncertain.
Answer:
In this case, option A represents MA, and option B represents CMA. While the model select MA as the final answer, its rationale indicates that it successfully received and processed the information from CMA. However, for some unknown reason, it still provides the MA as the final answer. This reasoning process is highly confusing. We suspect that the higher $R_{m}$ of LLaMA2.7B can be attributed to its weakness in reasoning ability. This finding raises an interesting question about the relationship between reasoning ability and memory strength of LLMs. We leave this for future work.
Appendix B Methodology and Experiment Details
B.1 CMA Generation for NQ dataset
@@
Table 6: Summary of Excluded Question Types in Memory Answer and Counter Answer Generation. The table lists question types that were excluded from processing due to either the difficulty in identifying a unified entity type (“how”, “what”, “which”) or poor question quality (“other”).
We generate CMA from MA with three steps: 1)identity question type, 2) determine entity type in MA to change, and 3) generate CMA with LLMs.
Identity Question Type: We first build a typing tree using rules to categorize questions. Figure 5 illustrates the typing tree, which consists of a two-layer structure. In the typing process, we first determine if a question begins with one of the following words: “what”, “when”, “where”, “which”, “who”, “why”, or “how”. If it does, the question is categorized accordingly; if not, it is classified as “other”. However, this approach can still group different types of questions together. To address this, we use a second layer to refine the typing by analyzing two specific words in the question. For example, the question shown in Figure 1 falls into the “how_many” category.
Determine Entity Type in MA to Change: After categorizing the questions, we determine the entity type in MA needs to be replaced. To achieve this, we give each type of question an entity type, and many questions can share the same entity type. For example, both “when” and “what year” ask for a time. So a time entity in MA should be substituted. The final set of entity types is summarized in Table 5. We do not process questions starting with “what”, “which” or “how” due to the lack of a unified entity type for these questions. Table 6 shows the statistics of the unprocessed questions.
B.2 Dataset Details
@@
LLaMA2.7B LLaMA2.70B ChatGPT GPT-4 LLaMA2.7B LLaMA2.70B ChatGPT GPT-4
Initial (# of Q) 1000 1000 1000 1000 1667 1667 1667 1667
Generate MA 1000 1000 1000 1000 1435 1392 1532 1539
Generate CMA 1000 1000 1000 1000 1152 1101 1189 1252
CMA filter 922 932 944 946 1060 1027 1110 1188
Direct Evidence 918 922 933 933 1042 1009 1079 1171
2 sentence indirect evidence 903 910 921 923 990 985 1038 1129
3 sentence indirect evidence 907 897 918 924 991 982 1041 1125
intersection of 2&3 sentence evidence 901 895 911 918 976 972 1025 1108
@@
For the popQA dataset, we use the dataset from Xie et al. (2024) by randomly selecting 1,000 questions from the data intersection of the conflicts generated by LLaMA2.7B, LLaMA2.70B, ChatGPT, and GPT-4. We use the MA and CMA from Xie et al. (2024) and only generate direct evidence and indirect evidence using our framework. For the NQ dataset, we use the test set from Longpre et al. (2021), which consists of 1,667 unique questions. The MA, CMA, and evidence are all generated with our framework. The dataset scale at each step is presented in Table 7.
B.3 Human Evaluation for Model Reliability
@@
B.4 Impact of Memory Strength with Different Evidence Styles
@@
To demonstrate the relationship between context-faithfulness and memory strength with other evidence styles, we categorize the questions in each dataset into four groups according to the memory strength intervals [-2, -1], (-1, -0.5], (-0.5, -0.25] and [-0.25, 0], The evidence styles are direct + paraphrase evidence with two sentences and indirect evidence with two sentences. Figures 6,7 show the result. The figures show that there is a clear positive correlation between memory strength and MA ratio for both evidence styles, which implies that this positive correlation between memory strength and MA ratio is general.
@@
Figure 6: Relationship between Memory Strength and Answer Ratio with Direct + Paraphrase Evidence with Two Sentences.
Figure 7: Relationship between Memory Strength and Answer Ratio with Indirect Evidence with Two Sentences.
Figure 8: Relationship between Memory Strength and Answer Ratio with Direct Evidence. The option order is Counter First.
Origin Question Paraphrased Questions Answers
Table 8: Origin Question, Paraphrased Questions, and Corresponding Answers

### v2

@@ 5 Conclusion
We investigate how context-faithful LLMs are to external evidence across two datasets, PopQA and NQ datasets, using LLaMA2-7B, LLaMA2-70B, ChatGPT, GPT-4, LLaMA3.2-3B and Claude3.5. Our findings highlight the critical role of memory strength in shaping LLM behavior. There is a clear positive correlation between memory strength and memory answer ratio. Furthermore, we demonstrate that paraphrasing significantly enhances the context faithfulness of LLMs across various models and datasets. These findings offer valuable insights for advancing research in retrieval-augmented generation and context-based LLM applications.
@@
Acknowledgments
The work was supported in part by the US National Science Foundation under grant NSF-CAREER 2237831. We also want to thank the anonymous reviewers for their helpful comments.
@@ Acknowledgments
Within this supplementary material, we elaborate on the following aspects:
  • A Methodology Details
  • B Additional Studies
  • C Prompts
Appendix A Methodology and Experiment Details
A.1 CMA Generation for NQ dataset
Departing from prior approaches Longpre et al. (2021) that use NER techniques Chen et al. (2021); Zhou et al. (2022); Li et al. (2025) to locate and substitute answer entities, we employ LLMs to automatically generate alternative entities. Our CMA construction process from MA consists of three steps: 1) identify question type, 2) determine entity type in MA to change, and 3) generate CMA with LLMs.
Identity Question Type: We first build a typing tree using rules to categorize questions. Figure 5 illustrates the typing tree, which consists of a two-layer structure. In the typing process, we first determine if a question begins with one of the following words: “what”, “when”, “where”, “which”, “who”, “why”, or “how”. However, this approach can still group different types of questions together. To address this, we use a second layer to refine the typing by analyzing two specific words in the question. For example, the question shown in Figure 2 falls into the “how_many” category. Table 4 shows the statistics of question types of the NQ dataset. Note that, we find 127 samples that are not questions in the process, so we list them as “other”.
Determine Entity Type in MA to Change: After categorizing the questions, we determine the entity type in MA needs to be replaced. To achieve this, we give each type of question an entity type, and many questions can share the same entity type. For example, both “when” and “what year” ask for a time. So a time entity in MA should be substituted. The final set of entity types is summarized in Table 5. We do not process questions starting with “what”, “which” or “how” due to the lack of a unified entity type for these questions. Table 6 shows the statistics of the unprocessed questions. These question types account for only 7.3% of the NQ dataset (with 2, 98, 22, and 0 instances for “how”, “what”, “which”, and “why”, respectively), and this issue does not occur in the PopQA dataset. Since the overall trends are consistent across both datasets, we believe the missing questions will not affect our conclusions.
@@ A.1 CMA Generation for NQ dataset
Table 6: Summary of Excluded Question Types in Memory Answer and Counter Answer Generation. The table lists question types that were excluded from processing due to either the difficulty in identifying a unified entity type (“how”, “what”, “which”) or not question (“other”).
A.2 Dataset Details
@@
LLaMA2-7B LLaMA2-70B LLaMA3.2-3B ChatGPT GPT-4 Claude3.5 LLaMA2-7B LLaMA2-70B LLaMA3.2-3B ChatGPT GPT-4 Claude3.5
Initial (# of Q) 1000 1000 1000 1000 1000 1000 1667 1667 1667 1667 1667 1667
Generate MA 1000 1000 1000 1000 1000 1000 1435 1392 1352 1532 1539 1482
Generate CMA 1000 1000 1000 1000 1000 1000 1152 1101 1140 1189 1252 1232
CMA filter 922 932 942 944 946 934 1060 1027 1123 1110 1188 1228
Direct Evidence 918 922 938 933 933 931 1042 1009 1002 1079 1171 1173
2 sentence indirect evidence 903 910 922 921 923 914 990 985 965 1038 1129 1116
3 sentence indirect evidence 907 897 925 918 924 920 991 982 980 1041 1125 1122
intersection of 2&3 sentence evidence 901 895 917 911 918 913 976 972 941 1025 1108 1059
@@
For the popQA dataset, we use the dataset from Xie et al. (2024) by randomly selecting 1,000 questions from the data intersection of the conflicts generated by LLaMA2-7B, LLaMA2-70B, LLaMA3.2-3B, ChatGPT, GPT-4 and Claude3.5. We use the MA and CMA from Xie et al. (2024) and only generate direct evidence and indirect evidence using our framework. For the NQ dataset, we use the test set from Longpre et al. (2021), which consists of 1,667 unique questions. The MA, CMA, and evidence are all generated with our framework. The dataset scale at each step is presented in Table 7.
A.3 Human Evaluation for Model Reliability
@@
Appendix B Additional Studies
B.1 Memory Strength on Different Datasets
We illustrate the distributions of memory strength on the popQA and NQ datasets for LLaMA2-7B, LLaMA2-70B, LLaMA3.2-3B, ChatGPT, GPT-4 and Claude3.5, respectively (shown in Figure 6). The results show that LLMs demonstrate stronger memory for the NQ dataset than the popQA dataset. For the NQ dataset, most questions fall within the bin of (0.25, 0]. Only a few questions fall within bins of weaker memory strength. In contrast, the popQA dataset has a greater number of questions in bins with weaker memory strength. This phenomenon is consistent across all six evaluated LLMs, indicating that the LLMs have more knowledge of the NQ dataset compared to the PopQA dataset. A possible explanation is that the NQ dataset covers more commonly discussed subjects than those in the PopQA dataset. These subjects may have been encountered more frequently during the training of the LLMs, making it easier for the models to recall the information and resulting in stronger memory strength.
B.2 Order of Options
Dataset Evidence Style S # LLaMA2-7B LLaMA2-70B LLaMA3.2-3B ChatGPT GPT-4 Claude3.5
$R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$ $R_{m}\downarrow$ $R_{c}\uparrow$ $R_{u}$
MA first
popQA Direct 1 0.44 99.56 0.0 1.08 98.7 0.22 7.69 85.15 7.16 3.32 94.75 1.93 13.29 84.57 2.14 0.65 83.31 16.04
Direct + Paraphrase 3 0.11 99.89 0.0 0.43 99.35 0.22 1.07 97.86 1.07 1.39 98.28 0.32 1.29 98.5 0.21 0.11 98.06 1.83
CMA first
Direct 1 59.37 40.41 0.22 5.75 93.27 0.98 18.8 69.12 12.07 6.22 91.96 1.82 21.44 75.24 3.32 0.43 76.96 22.61
Direct + Paraphrase 3 17.49 82.51 0.0 1.74 98.05 0.22 14.64 79.92 5.45 1.82 97.75 0.43 2.79 96.78 0.43 0.21 96.88 2.91
MA first
NQ Direct 1 7.2 92.8 0.0 3.07 96.93 0.0 26.41 59.88 13.71 19.46 75.16 5.38 50.04 47.99 1.96 1.96 56.4 41.64
Direct + Paraphrase 3 3.26 96.74 0.0 1.19 98.61 0.2 9.38 83.27 7.36 9.55 86.75 3.71 11.27 87.28 1.45 0.26 93.09 6.65
CMA first
Direct 1 22.26 77.73 0.0 19.13 80.38 0.5 41.33 42.44 16.23 34.48 61.82 3.71 49.19 47.99 2.82 4.78 39.85 55.38
Direct + Parapharse 3 4.8 95.11 0.1 8.72 90.39 0.89 24.19 65.62 10.18 18.63 78.96 2.41 17.76 80.02 2.22 1.28 78.84 19.88
Table 8: Results of LLM Receptiveness to Different Evidence Styles Across popQA and NQ Datasets. The table presents the MA ratio ($R_{m}$), CMA ratio ($R_{c}$), and UCT ratio ($R_{u}$) for Direct Evidence and Direct + Paraphrase Evidence with CMA first and MA first scenarios. All the ratios are in %.
Figure 7: Impact of Option orders on Memory and Counter Ratios Across NQ and popQA Datasets. Either the memory answer ("mem first") or the counter answer ("ctr first") is introduced first to six models.
Across all six models (LLaMA2-7B, LLaMA2-70B, LLaMA3.2-3B, ChatGPT, GPT-4 and Claude3.5), we observe a consistent trend: MA ratio ($R_{m}$) under “CMA first” is significantly higher than that under “MA first”. Evaluations under “CMA first” demonstrate that LLMs are less context-faithful.
To further demonstrate the effect of the order of options on $R_{m}$, we compare the performance of experiments with “CMA first” and “MA first” under two evidence styles: Direct Evidence with one sentence and Direct + Paraphrase with three sentences. The results are presented in Table 8. The results show that, for different evidence styles, $R_{m}$ is higher in the “CMA first” compared to the “MA first”. Comparing the results under the “CMA first”, the $R_{m}$ of Direct + Paraphrase with three sentences is significantly lower than that of Direct Evidence with one sentence. This demonstrates that paraphrasing direct evidence is an effective method for decreasing $R_{m}$. Our conclusion remains unchanged.
B.3 Entity Type
To investigate how entity type influences context faithfulness, we categorize NQ questions into three types based on their expected answer entities: PER (person), LOC (location), and TIM (time), following the classification summarized in Table 3. Specifically:
  • PER includes questions such as “who”, “what_name”, “who_writes”, “who_sings”, “who_wins”, and “who_plays”.
  • LOC includes “where” and “which_country”.
  • TIM includes “when”, “what_year”, and “which_year”.
We then analyze the MA ratio ($R_{m}$) across different entity types using ChatGPT and LLaMA2-7B as representative models. For each category, we compute the question count, average memory strength, and MA ratio. Results are shown in Table 9. We observe the following patterns:
LLM Entity Count Mem. Strength MA Ratio (%)
ChatGPT PER 664 -0.4467 24.54
LOC 107 -0.5033 19.62
TIM 217 -0.3244 9.21
LLaMA2-7B PER 680 -0.6149 8.52
LOC 95 -0.4344 9.47
TIM 174 -0.4359 1.15
Table 9: Impact of different entity types (PER: Person, LOC: Location, TIM: Time) on memory strength and MA ratio for ChatGPT and LLaMA2-7B.
LLM low mid-low mid-high high
ChatGPT 4.76 9.09 9.09 10.08
LLaMA2-7B 0.00 2.13 0.00 2.53
Table 10: MA ratios (%) on TIM-type questions across different memory strength levels.
  • For PER and LOC, which involve textual entities, higher memory strength corresponds to higher MA ratios. This is consistent with the trend reported in Section 4.2.1, where stronger memory increases the likelihood of memory-aligned responses.
  • For TIM, which involves numerical entities, the MA ratio is substantially lower than for textual entities. This finding aligns with observations made in Longpre et al. (2021).
  • We further divided TIM into four groups based on memory strength levels: low, mid-low, mid-high, and high. The results shown in Table 10 and these results demonstrate a positive correlation between memory strength and MA ratio, which aligns with the conclusion presented in Section 4.2.1.
These results suggest that entity type, particularly the distinction between textual and numerical entities, affects context faithfulness, and that memory strength remains a key influencing factor across different entity types.
B.4 Case Study
Poor reasoning ability may be one of the reasons that LLMs provide MA. In our case study, we prompt LLMs to provide a rationale alongside the answer. We discover that some LLMs seem to exhibit reasoning errors. Below is an example that shows this phenomenon.
Question: who has the most big ten championships in football
Options:
C: Uncertain.
Answer:
In this case, option A represents MA, and option B represents CMA. While the model select MA as the final answer, its rationale indicates that it successfully received and processed the information from CMA. However, for some unknown reason, it still provides the MA as the final answer. This reasoning process is highly confusing. We suspect that the cause of MA can be attributed to LLMs’ weakness in reasoning ability. This finding raises an interesting question about the relationship between reasoning ability and memory strength of LLMs. We leave this for future work.
B.5 Impact of Memory Strength with Different Evidence Styles
@@
To demonstrate the relationship between context-faithfulness and memory strength with other evidence styles, we categorize the questions in each dataset into four groups according to the memory strength intervals [-2, -1], (-1, -0.5], (-0.5, -0.25] and [-0.25, 0], The evidence styles are direct + paraphrase evidence with two sentences and indirect evidence with two sentences. Figures 8,9 show the result. The figures show that there is a clear positive correlation between memory strength and MA ratio for both evidence styles, which implies that this positive correlation between memory strength and MA ratio is general.
@@
Figure 8: Relationship between Memory Strength and Different Answers’ Ratios with Direct + Paraphrase Evidence with Two Sentences.
Figure 9: Relationship between Memory Strength and Different Answers’ Ratios with Indirect Evidence with Two Sentences.
Figure 10: Relationship between Memory Strength and Different Answers’ Ratios with Direct Evidence. The option order is Counter First.